A jury pool consists of 35 people, 18 men and 17 women. Compute the probability that a randomly selected jury of 12 people is all male.

Answer by Edwin McCravy(20055)

C(18,12)/C(35,12) = 18564/834451800

You can reduce that fraction by dividing top and bottom by the top:

1/44950 approximately = 0.00002...

Edwin